# HyperSHAP: Shapley Values and Interactions for Hyperparameter Importance

### v1

@@ E.3 Additional Results for Explaining the SMAC Surrogate During Optimization
In Figure 12, in addition to the MI interaction graphs, we summarize explanations with the help of second order FSII, which fairly distributes higher-order interactions to the lower orders, here order one and two. We find that with FSII we can distill the relevant parts of the MIs quite clearly.
@@
Figure 12: HyperSHAP Tunability explanations for the surrogate model used in SMAC at different time intervals (1%, 5%, 25%, 100%) of the optimization procedure for dataset 3945 of lcbench (Zimmer et al., 2021). Over time the model becomes less uncertain about which hyperparameters are important to achieve a high predictive performance. Bottom: Interaction graphs for Moebius Interactions (MI) show all pure main effects and interactions. Top: Higher-order interactions are summarized to main effects and second-order interactions, summarizing the game properly already at early stages when the MI still shows a comparably large number of higher-order interactions.
@@ E.4 Additional Interaction Visualizations
In Figures 13, 14, 15 and 16, we show more interaction graphs for the different benchmarks we evaluated HyperSHAP on. This includes PD1 (cf. Figure 13, JAHS-Bench-201 (cf. Figure 14), lcbench (cf. Figure 15), and rbvs_ranger (cf. Figure 16). We find that with HyperSHAP we can elicit interesting interaction structures for the tuning of transformers and neural architectures in more general. Surprisingly, there can be comparably low interaction between hyperparameters steering the learning behavior and hyperparameters controlling the neural architecture, as seen for CIFAR10. However, for the other two datasets, the higher degree of interaction between the learner’s hyperparameters and those of the architecture better meets intuition and expectation.
In Figure 17, we compare the Sensitivity to the Tunability game for dataset ID 7593 of lcbench on three different levels: Moebius interactions showing all pure effects, Shapley interactions, summarizing higher-order interactions to main effects and interactions of order two and Shapley values representing the entire game solely in terms of main effects. What we can observe is that Tunability and Sensitivity yield quite different explanations as Sensitivity does not blend an optimized hyperparameter configuration with the default hyperparameter configuration for evaluating the value function for a given coalition but takes the variance. Taking the variance apparently results in more pronounced interactivity structures as the performance is no longer contrasted to the default configuration.

### v2

@@ E.3 Additional Results for Explaining the SMAC Surrogate During Optimization
In Fig. 12, in addition to the MI interaction graphs, we summarize explanations with the help of second order FSII, which fairly distributes higher-order interactions to the lower orders, here order one and two. We find that with FSII we can distill the relevant parts of the MIs quite clearly.
@@
Figure 12: HyperSHAP Tunability explanations for the surrogate model used in SMAC at different time intervals (1%, 5%, 25%, 100%) of the optimization procedure for dataset 3945 of lcbench (zimmer-tpami21a). Over time the model becomes less uncertain about which hyperparameters are important to achieve a high predictive performance. Bottom: Interaction graphs for Moebius Interactions (MI) show all pure main effects and interactions. Top: Higher-order interactions are summarized to main effects and second-order interactions, summarizing the game properly already at early stages when the MI still shows a comparably large number of higher-order interactions.
@@ E.4 Additional Interaction Visualizations
In Figs. 13, 14, 15 and 16, we show more interaction graphs for the different benchmarks we evaluated HyperSHAP on. This includes PD1 (cf. Fig. 13, JAHS-Bench-201 (cf. Fig. 14), lcbench (cf. Fig. 15), and rbvs_ranger (cf. Fig. 16). We find that with HyperSHAP we can elicit interesting interaction structures for the tuning of transformers and neural architectures in more general. Surprisingly, there can be comparably low interaction between hyperparameters steering the learning behavior and hyperparameters controlling the neural architecture, as seen for CIFAR10. However, for the other two datasets, the higher degree of interaction between the learner’s hyperparameters and those of the architecture better meets intuition and expectation.
In Fig. 17, we compare the Sensitivity to the Tunability game for dataset ID 7593 of lcbench on three different levels: Moebius interactions showing all pure effects, Shapley interactions, summarizing higher-order interactions to main effects and interactions of order two and Shapley values representing the entire game solely in terms of main effects. What we can observe is that Tunability and Sensitivity yield quite different explanations as Sensitivity does not blend an optimized hyperparameter configuration with the default hyperparameter configuration for evaluating the value function for a given coalition but takes the variance. Taking the variance apparently results in more pronounced interactivity structures as the performance is no longer contrasted to the default configuration.
@@
Ethical Statement
We believe that HyperSHAP can support practitioners in understanding the role of hyperparameters in model performance, quantify tunability, making model-based optimization more transparent, and uncover biases in optimizers. This is particularly valuable in resource-constrained settings, where informed decisions about which hyperparameters to tune can render HPO more efficient.